# BPDec: Unveiling the Potential of Masked Language Modeling Decoder in BERT Model Pretraining

### v3

@@ Abstract
BERT (Bidirectional Encoder Representations from Transformers) has revolutionized the field of natural language processing through its exceptional performance on numerous tasks. Yet, the majority of researchers have mainly concentrated on enhancements related to the model structure, such as relative position embedding and more efficient attention mechanisms. Others have delved into pretraining tricks associated with Masked Language Modeling, including whole word masking. DeBERTa introduced an enhanced decoder adapted for BERT’s encoder model for pretraining, proving to be highly effective. We argue that the design and research around enhanced masked language modeling decoders have been underappreciated. In this paper, we propose several designs of enhanced decoders and introduce BPDec (BERT Pretraining Decoder), a novel method for modeling training. Typically, a pretrained BERT model is fine-tuned for specific Natural Language Understanding (NLU) tasks. In our approach, we utilize the original BERT model as the encoder, making only changes to the decoder without altering the encoder. This approach does not necessitate extensive modifications to the encoder architecture and can be seamlessly integrated into existing fine-tuning pipelines and services, offering an efficient and effective enhancement strategy. Compared to other methods, while we also incur a moderate training cost for the decoder during the pretraining process, our approach does not introduce additional training costs during the fine-tuning phase. We test multiple enhanced decoder structures after pretraining and evaluate their performance on the GLUE tasks and SQuAD tasks. Our results demonstrate that BPDec, having only undergone subtle refinements to the model structure during pretraining, significantly enhances model performance without escalating the finetuning cost, inference time and serving budget.
@@ 1 Introduction
The journey of deep learning is marked by significant innovations with propelling advancements that redefine our understanding and capabilities in solving complex problems [1, 2, 3]. ResNet revolutionized a wide range of computer vision tasks with its deep residual learning framework, enabling the training of deeper networks by addressing the vanishing gradient problem [2]. This breakthrough laid the groundwork for subsequent works in model architectures. The transformer model shifted the paradigm, establishing a new standard in handling sequential data [4]. Building on these foundations, BERT model marked another milestone in natural language processing. By pretraining on a large corpus and adopting a bidirectional approach, BERT achieved unprecedented performance across a wide range of language tasks. These contributions collectively have not only advanced the field of deep learning but also have set the stage for our current exploration into refining the BERT architecture through a pretraining and finetuning approach.
Following the groundbreaking inception of the Transformer architecture, there has been a continuous evolution in model design and methodology, further enhancing its capabilities and performance in diverse tasks. TransformerXL extended the Transformer’s capability to process longer sequences, overcoming the limitations of the standard architecture in handling extended context [5]. This was achieved through the innovative use of relative positional encoding, allowing the model to capture dependencies over extended sequences, far beyond the absolute positional encoding that was applied in Transformer and BERT models. This enhancement not only addressed the limitations in handling extended context but also paved the way for more dynamic and contextually aware representations in sequence modeling. Complementing this, XLNet integrated the strengths of both autoregressive and autoencoding techniques [6]. It further leveraged the concept of relative positional encoding, enhancing the model’s ability to understand and predict the relationships between elements in a sequence. This approach enabled XLNet to surpass BERT in numerous benchmarks. The Funnel-Transformer, another significant innovation, introduces a methodology that gradually compresses the sequence of hidden states into a shorter format, effectively reducing computation costs without compromising the model’s performance across a variety of NLP tasks [7]. This trend underscores a growing emphasis in the field not only on achieving high accuracy but also on improving computational efficiency.
@@
The original Transformer architecture has a separation of the encoder and decoder components, each serving a unique function in the model’s processing of sequences. BERT focused exclusively on the encoder part of this architecture. By harnessing the power of the encoder, BERT excelled in creating deep, bidirectional representations with Masked Language Modeling (MLM), but it lose the sequential generation capabilities of the decoder. On the other hand, T5 (Text-to-Text Transfer Transformer) took a holistic approach by further developing both the encoder and decoder components [8]. T5 not only embraced the full encoder-decoder structure but also innovated by treating every NLP task as a text-to-text problem and using a multi-tasking pretraining strategy. This model incorporated varied methods such as MLM with different lengths and formats, thereby expanding the versatility and applicability of the Transformer architecture. Such developments in models like T5 demonstrate the continuous evolution and adaptation of the Transformer framework, exploring new potentials in both its encoder and decoder capabilities. These advancements reflect that, in addition to improvements in model architecture, changes in strategies during the pretraining phase are also very important.
@@
Not surprisingly, alongside these architectural advancements, there has been a significant emphasis on refining strategies and methods in optimization, hyperparameter tuning, and model pretraining. For instance, RoBERTa, an iteration of BERT, has unveiled its potential by utilizing more extensive and diverse datasets, dynamic masking in Masked Language Modeling (MLM), and a refined set of hyperparameters [9]. This work underlines the importance of not just model design, but also the quality and variety of data and training techniques. Furthermore, some studies have delved into the optimal ratio of masking in MLM, a key factor that significantly influences the efficiency and effectiveness of the pretraining process [10]. Moreover, the topic of how to mask tokens has also been widely discussed [11]. Another intriguing area of research is the adaptability and flexibility of training structures. Innovations like BERT-of-Theseus demonstrate how a compressing strategy can lead to more efficient models without compromising performance [12]. These strategies involve progressively replacing parts of the network with more efficient modules, thereby refining the model’s structure and training process. Such developments suggest a growing recognition of the intricate interplay between various aspects of model training, extending beyond mere architectural changes to include strategic modifications in data handling, training methodologies, and structural efficiency.
DeBERTa represents a significant milestone in the ongoing evolution of language models, introducing several novel methods that have achieved state-of-the-art performance [13]. The authors innovatively introduced the disentangled attention mechanism represents a significant deviation from vanilla multi-head attention. However, compared to a BERT model with a same number of layers and similar amount of parameters, DeBERTa incurs a higher computational cost for both training and serving. Furthermore, for developers considering transitioning from BERT to DeBERTa, there is an additional development cost need to be considered, which may not always be feasible or cost-effective for every application or organization.
@@
However, DeBERTa’s introduction of a pretraining-only module called the Enhanced Mask Decoder, which is essentially two additional layers discarded during fine-tuning, intrigued us more than previous architectural modifications. We were drawn to explore whether this decoder module could be generalized and optimized for use with the vanilla BERT model, and whether it could unlock encoder’s ability to match or even surpass the performance of more computationally expensive models. This line of inquiry forms the basis of our research.
@@
In this paper, we propose BPDec (BERT Pretraining Decoder), a novel architecture that harnesses the potential of the MLM (Masked Language Modeling). Our primary contributions are as follows:
1. Introduction of BPDec: Its encoder is identical to the vanilla BERT, but it includes a redesigned MLM decoder called BPDec. The main modifications are:
@@
  • Adding multiple transformer block layers after the encoder to function as a decoder.
  • In the decoder, we remove the restriction present in the encoder’s attention mechanism, which prevents attending to masked positions. Instead, we randomly and gradually relax some of the masked positions throughout the decoder layers.
@@
2. Guidance on Hyperparameter Tuning for BPDec: To further assist readers, we have incorporated a dedicated section in our paper to guide the tuning process of the MLM decoder’s hyperparameters.
To assess the performance of our proposed method, we conducted a series of rigorous evaluations across a variety of tasks. These tasks include, but are not limited to, MNLI (Multi-Genre Natural Language Inference), SQuAD (Stanford Question Answering Dataset), RACE (Reading Comprehension from Examinations), among others. In addition, an extensive ablation study was performed to validate the effectiveness of each modification we made. This study methodically deconstructed the model to isolate and evaluate the impact of individual changes, thereby providing empirical evidence of their contribution to the model’s overall performance. Furthermore, comparative analysis were conducted against baseline models, including the original BERT and other state-of-the-art architectures. These comparisons not only highlight BPDec’s advancements but also offer a comprehensive understanding of its positioning in the current NLP landscape.
@@ 2.1 Masked Language Modeling
Masked Language Modeling (MLM) has become a fundamental technique in the field of natural language processing, particularly since the introduction of the BERT model [13, 9]. BERT represents a significant shift in model architecture by employing an encoder-only framework [1]. For autoregressive language modeling, as seen in models like GPT, the objective differs in its sequential approach to predicting each token. In autoregressive language modeling, each token $X_{i}$ is predicted based on the sequence of preceding tokens. Mathematically, this can be defined as:
$\max_{\theta}\log p_{\theta}(X|\tilde{X})=\max_{\theta}\sum_{i=1}^{N}\log p_{\theta}(X_{i}|X_{1},X_{2},...,X_{i-1})$ (2)
The goal is to optimize the model parameters $\theta$ such that the probability of each token $X_{i}$, given its preceding context, is maximized. The encoder-only design of BERT enables bidirectional context understanding and full-text attention, unlike autoregressive decoders that are limited to unidirectional context. BERT’s training methodology evolves to masked language modeling (MLM). This approach can be mathematically represented as follows:
@@
Here, $X_{masked}$ represents the set of tokens that have been masked, and $X_{unmasked}$ is the set of tokens that remain visible to the model. The model parameters $\theta$ are optimized to maximize the log probability of correctly predicting the masked tokens given the context of unmasked tokens. BERT’s MLM pretraining randomly masks 15% of the tokens in each sequence. Within this subset, 80% of the tokens are substituted with a [MASK] token, 10% remain as is, and the remaining 10% are replaced with a random token from the vocabulary. The model is trained to predict these tokens based on the context. This training method allows BERT to effectively grasp the bidirectional context, making it highly effective across a wide range of NLP tasks after finetuning.
@@ 2.2 Finetuning
There are variations in the finetuning process. Some approaches suggest the possibility of dropping certain layers during fine-tuning for efficiency [14]. This efficiency-focused approaches suggest the feasibility of omitting pretrained layers during fine-tuning. Notable findings in this area highlight that:
@@
2.3 Enhanced Masked Language Modeling Decoder in DeBERTa
DeBERTa introduces an innovative approach known as the Enhanced Masked Language Modeling Decoder (EMD), which is mentioned above. Moreover, while BERT incorporates absolute positional encodings at the input layer, DeBERTa adopts a different strategy. In DeBERTa, absolute positions are incorporated after all Transformer layers, just before the softmax layer dedicated to masked token prediction. This design choice allows the model to capture the relative positions within all Transformer layers and utilize absolute positions as supplementary information during the decoding phase. However, this disentangled attention mechanism notably introduces additional parameters and computational overhead compared to the original attention mechanism. Therefore, a direct comparison between DeBERTa and BERT with same number of layers might not be entirely equitable due to these architectural differences. As we will elaborate in the next section, our approach ensures fairness in comparison. The only difference lies in the pretraining phase compared to the original BERT, which is often just a one-time cost, whereas in practical applications, finetuning and inference are the recurring stages where computational efficiency matters more.
@@ 3 Methods
Considering the structural modifications of DeBERTa and their impact on the overall computational cost of the model, we find the concept of Enhanced Mask Decoders (EMD) to be particularly intriguing. The MLM decoder offers a solution that it is utilized exclusively during the pretraining phase, thereby avoiding additional computational burdens during the finetuning and serving. This aspect caught our attention and led us to explore the potential it holds for BERT pretraining. While the primary intent of introducing the Enhanced Mask Decoder (EMD) in the DeBERTa model, as highlighted by the authors, was to integrate absolute position information before the output layer. However, insights from other studies suggest an alternative perspective [14]. The ability to understand the input text and the ability to distill language into specific embeddings already well-established in the earlier layers, which indicates that using later layer during BERT finetuning might be unnecessary.
In related work, we also discussed T5 and BERT-of-Theseus demonstrate that variation between pretraining and fine-tuning architectures can yield stronger performance [8, 12]. Building upon these insights, we hypothesize that carefully designed differences in these architectures can not only avoid performance degradation but may even enhance some part of the model. This leads us to believe that, with further research and meticulous design, the original BERT model itself has untapped potential for significant improvement. Motivated by these insights and our analysis, we developed the BERT Pretraining MLM Decoder, termed BPDec, a novel architecture that serves as an innovative paradigm for BERT pretraining.
@@ 3.1 MLM Decoder
Our approach with BPDec preserves the core encoder structure of the original BERT while augmenting it with a specially designed decoder, exclusively active during pretraining. This additional decoder is crafted to enhance the BERT encoder’s ability in understanding and processing of language, drawing upon the strengths of the EMD approach. By incorporating this decoder, we aim to improve the overall performance of the BERT encoder without imposing significant computational costs during the subsequent phases of finetuning and deployment. The BPDec design is strategically aligned with the objectives of pretraining. This decoder is adeptly positioned right after the BERT encoder, and just before the final dense layer and softmax layer dedicated to masked token prediction. It focuses on refining the model’s ability to predict masked tokens, thereby expanding the depth and versatility of the encoded representations.
3.2 Gradual Relaxation of Attention Mask
@@
Figure 2: Examples of gradual relaxation of attention mask. (a) Encoder layers are unaffected. (b) Randomly unmask a portion of the masked positions. (c) In the final layer of the BPDec, Unmask the rest of the attention mask and fully disable the attention mask in the .
@@
However, in our MLM Decoder, we try to remove the restriction in the decoder layers. This modification is rooted in the understanding that as the layers in the model become deeper, the output features become increasingly abstract and, conceivably, become sufficiently enough to guess the masked or noised information. Such information, though not fully revealed, can provide valuable context to the model. It encourages the model to formulate hypotheses and fosters a deeper alignment and understanding of the input based on a mix of noise and these formulated assumptions. However, this also brings about a drawback, which is that the model’s sudden exposure to this hypothesized information about the masked or noised input, and the noise it brings, may have adverse effects. To mitigate these risks and enhance the effects, we made two optimizations shown in Figure 2: 1) Fully disabling the attention mask only in the last several decoder layers to prevent the early involvement of hypothesized information, and 2) Progressively unmasking the attention mask across the decoder layers, starting with a randomized partial unmasking followed by a complete unmasking. This approach not only enriches the model’s interpretative capacity but also allows for a more gradual training process, leading to enhanced overall performance as shown in the experiments.
@@ 3.3 Random Mix of Encoder and Decoder Outputs
In our approach, we introduce a modest blending of the encoder’s output into the decoder’s output as shown in Figure 3. This technique serves a dual purpose. Firstly, it encourages the MLM task to not depend exclusively on the representations formed in the deeper layers of the decoder. Instead, it also takes into account the more direct and less processed information from the encoder stage. Drawing inspiration from the NEFTUNE method, which highlights the potential performance benefits of introducing controlled randomness during language model training, we adopt a similar principle in our approach [15]. Additionally, echoing the compression method utilized in BERT-of-Theseus, the authors recognize that random layer replacement is an effective approach to model compression [12]. In our case, randomly skipping decoder layers can be seen as a form of compression as well. Our experimental results reveal that using 80% of the decoder output mixed with 20% of the encoder output can significantly improve performance on downstream tasks.
@@ 4 Results
In order to examine the benefits of BPDec, we conduct experiments on various NLP tasks with a few models with as baselines.
@@ 4.1 Pretraining
Pretraining was conducted on both base and large model sizes. Additionally, we introduced a variant of the base model (base-h256) to further demonstrate the generalizability of our method on smaller models. The encoder architecture, pretraining data, pretraining tasks, and mask settings are identical with the original BERT models. Our BPDec lies in the addition of a decoder following the encoder. Details about the parameters of this decoder can be found in Table 1. This strategic augmentation aims to enhance the model’s capabilities while maintaining the foundational architecture of BERT encoder, and it also ensures a high degree of fairness in our comparative experiments.
@@
Moreover, it is important to note that although the DeBERTa model was trained for the same number of steps as other models, its disentangled attention mechanism introduced a significant overhead comparing with other baseline models and with our BERT+BPDec models. This added complexity affected both the pretraining and fine-tuning phases, but ours BERT+BPDec models only have extra ovearhead during pretraining. The implications of this increased overhead are critical, especially when considering the balance between model performance and computational efficiency. Our analysis aims to provide insights into how these different architectures impact training and computation efficiency, while also maintaining or enhancing language processing capabilities.

### v4

@@ Abstract
BERT has revolutionized the field of natural language processing through its exceptional performance on numerous tasks. Yet, the majority of researchers have mainly concentrated on enhancements related to the model structure and pretraining, such as relative position embedding, more efficient attention mechanisms, and Masked Language Modeling, including whole word masking. In this paper, we introduce BPDec (BERT Pretraining Decoder), a novel method for modeling pretraining. BPDec is enhanced with an addition of multiple transformer blocks as a decoder, the introduction of Gradual Unmasking Attention (GUA), and the integration of output randomness for a more dynamic final output. We utilize the original BERT model as the encoder so this approach does not necessitate extensive modifications to the encoder architecture and can be seamlessly integrated into existing fine-tuning pipelines and services, offering an efficient and effective enhancement strategy with no additional training cost during the fine-tuning phase. We test multiple enhanced decoder structures after pretraining and evaluate their performance on the GLUE tasks and SQuAD tasks. Our results demonstrate that BPDec, having only undergone subtle refinements to the model structure during pretraining, significantly enhances model performance without escalating the finetuning cost, inference time and serving budget.
Keywords:
Natural Language Processing Language Modeling Foundation Model Pretraining.
@@ 1 Introduction
Following the groundbreaking inception of the Transformer architecture, there has been a continuous evolution in model design and methodology, further enhancing its capabilities and performance in diverse tasks. The original Transformer architecture has a separation of the encoder and decoder components, each serving a unique function in the model’s processing of sequences. BERT focused exclusively on the encoder part of this architecture. By harnessing the power of the encoder, BERT excelled in creating deep, bidirectional representations with Masked Language Modeling (MLM), but it lost the sequential generation capabilities of the decoder. On the other hand, T5 (Text-to-Text Transfer Transformer) took a holistic approach by further developing both the encoder and decoder components [1]. T5 not only embraced the full encoder-decoder structure but also innovated by treating every NLP task as a text-to-text problem and using a multi-tasking pretraining strategy. Besides, XLNet integrated the strengths of both autoregressive and autoencoding techniques [2, 3]. It further leveraged the concept of relative positional encoding, enhancing the model’s ability to understand and predict the relationships between elements in a sequence.
@@
There has been a significant emphasis on refining strategies and methods in optimization, hyperparameter tuning, and model pretraining. For instance, RoBERTa, an iteration of BERT, has unveiled its potential by utilizing more extensive and diverse datasets, dynamic masking in Masked Language Modeling (MLM), and a refined set of hyperparameters [4]. This work underlines the importance of not just model design, but also the quality and variety of data and training techniques. Furthermore, some studies have delved into the masking strategy, a key factor that significantly influences the efficiency and effectiveness of the pretraining process [5, 6]. Moreover, BERT-of-Theseus demonstrate how a compressing pretraining method can lead to more efficient models without compromising performance [7].
@@
DeBERTa represents a significant milestone in the ongoing evolution of language models, introducing the disentangled attention mechanism represents a significant deviation from vanilla multi-head attention [8]. However, compared to a BERT model with the same number of layers and similar amount of parameters, DeBERTa incurs a higher computational cost for both training and serving. Furthermore, for developers considering transitioning from BERT to DeBERTa, there is an additional development cost need to be considered, which may not always be feasible or cost-effective for every application or organization. However, DeBERTa’s introduction of a pretraining-only module called the Enhanced Mask Decoder, which is essentially two additional layers discarded during fine-tuning, intrigued us more than previous architectural modifications. We were drawn to explore whether this decoder module could be generalized and optimized for use with the vanilla BERT model, and whether it could unlock encoder’s ability to match or even surpass the performance of more computationally expensive models. This line of inquiry forms the basis of our research.
@@
In this paper, we propose BPDec (BERT Pretraining Decoder), a novel architecture that harnesses the potential of the MLM (Masked Language Modeling). We introduce a novel architecture BPDec. Its encoder is identical to the vanilla BERT, but it includes a redesigned MLM decoder called BPDec. The main modifications are:
@@
  • Adding multiple transformer layers after the encoder to function as a decoder.
  • We proposed a new attention mechanism called Gradual Unmasking Attention (GUA) for the pretraining decoder. Conventionally, the attention mechanism for BERT is applied with a mask that prevents attending to masked positions. Instead, in Gradual Unmasking Attention, we remove this restriction and allows the model to attend to the embeddings at masked positions randomly and gradually.
@@
To assess the performance of our proposed method, we conducted a series of rigorous evaluations and ablation studies across a variety of NLP tasks.
@@ 2.1 Masked Language Modeling
Masked Language Modeling (MLM) has become a fundamental technique in the field of natural language processing, particularly since the introduction of the BERT model [8, 4]. BERT represents a significant shift in model architecture by employing an encoder-only framework [9]. Here, $X_{masked}$ represents the set of tokens that have been masked, and $X_{unmasked}$ is the set of tokens that remain visible to the model. The model parameters $\theta$ are optimized to maximize the log probability of correctly predicting the masked tokens given the context of unmasked tokens. BERT’s MLM pretraining randomly masks 15% of the tokens in each sequence. Within this subset, 80% of the tokens are substituted with a [MASK] token, 10% remain as is, and the remaining 10% are replaced with a random token from the vocabulary. The model is trained to predict these tokens based on the context.
@@ 2.2 Finetuning
Some approaches suggest the possibility of dropping certain layers during fine-tuning for efficiency [10]. Notable findings in this area highlight that:
@@ 3 Methods
Considering the structural modifications of DeBERTa and their impact on the overall computational cost of the model, we find the concept of Enhanced Mask Decoders (EMD) to be particularly intriguing. The MLM decoder offers a module that is utilized exclusively during the pretraining phase, thereby avoiding additional computational burdens during the finetuning and serving. The primary intent of introducing the Enhanced Mask Decoder (EMD) in the DeBERTa model, was to integrate absolute position information before the output layer. However, insights from other studies suggest an alternative perspective [10]. The ability to understand the input text and the ability to distill language into specific embeddings are already well-established in the earlier layers, which indicates that using later layer during BERT finetuning might be unnecessary.
In related works, we observe that models like T5 and BERT-of-Theseus demonstrate that variations between pretraining and finetuning architectures do not degrade the model’s capabilities, particularly in the encoder part. On the contrary, these variations can potentially enhance performance [1, 7]. Building upon these insights, we hypothesize that carefully designed differences in these architectures can not only avoid performance degradation but may even enhance some part of the model. This leads us to believe that, with further research and meticulous design, the original BERT model itself has untapped potential for significant improvement. Motivated by these insights and our analysis, we propose the BERT Pretraining MLM Decoder, termed BPDec, a novel architecture that serves as an innovative paradigm for BERT pretraining.
@@ 3.1 MLM Decoder
Our approach with BPDec maintains the fundamental encoder architecture of the original BERT, but supplements it with a specially designed decoder, which is used exclusively during pretraining. This additional decoder, consisting of transformer blocks similar to those in the encoder, is crafted to enhance the BERT encoder’s ability in understanding languages. By incorporating this pretraining-only decoder, we aim to improve the overall performance of the BERT encoder without imposing significant computational costs during the subsequent phases of finetuning and deployment. The BPDec is strategically positioned after the encoder and just before the output layers for masked token prediction. It specifically sharpens the model’s ability to predict masked tokens, enriching the encoded representations with greater depth.
3.2 Gradual Unmasking Attention (GUA)
@@
Figure 2: Examples of gradual relaxation of attention mask. (a) Encoder layers are unaffected. (b) Randomly unmask a portion of the masked positions. (c) In the final layer of the BPDec, we unmask the rest of the attention mask and fully disable the attention mask.
@@
In the development of our new pretraining strategy, we introduce the Gradual Unmasking Attention (GUA) within our BPDec. This pioneering modification challenges the conventional constraints of the attention mechanism used in BERT model pretraining. Our approach is based on the insight that as the model’s layers progress deeper, the resulting output features become more abstract. We hypothesize that these abstracted features at masked positions are sufficiently accurate and detailed to serve as effective hypothesized contexts. GUA progressively unveils the attention mechanism to those hypothesized contexts of the masked inputs, fostering a more nuanced and robust formulation of language understanding. This method begins with partially unmasking the attention mask in the initial layers of the decoder. For instance, in a decoder with two transformer blocks as shown in Figure 2, the Gradual Unmasking Attention (GUA) in the first decoder layer would randomly unmask 50% of the masked positions. This introduces a controlled amount of noise and hypothesized contexts. In the second GUA, the remaining 50% of the positions are unveiled, allowing the model to integrate the full context. This phased approach of exposure ensures that the model gradually adapts to the changes in the inputs without being overwhelmed by sudden floods of unmasked data. More importantly, allowing the model to progressively form a complete picture of the input data significantly improves its interpretative capabilities, enhancing its ability to generate accurate and contextually relevant predictions.
@@ 3.3 Random Mix of Encoder and Decoder Outputs
In our approach, we introduce a modest blending of the encoder’s output into the decoder’s output as shown in Figure 3. This technique serves a dual purpose. Firstly, it encourages the MLM task to not depend exclusively on the representations formed in the deeper layers of the decoder. Instead, it also takes into account the more direct and less processed information from the encoder stage. Drawing inspiration from the NEFTUNE method, which highlights the potential performance benefits of introducing controlled randomness during language model training. We adopt a similar principle in our approach [11]. Additionally, echoing the compression method utilized in BERT-of-Theseus, the authors recognize that random layer replacement is an effective approach to model compression [7]. In our case, randomly skipping decoder layers can be seen as a form of compression as well. Our experimental results reveal that using 80% of the decoder output mixed with 20% of the encoder output can significantly improve performance on downstream tasks.
@@ 4.1 Pretraining
Pretraining was conducted on both base and large model sizes. Additionally, we introduced a variant of the base model (base-h256) to further demonstrate the generalizability of our method on smaller models. The encoder architecture, pretraining data, pretraining tasks, and mask settings are identical with the original BERT models. Our BPDec lies in the addition of a decoder following the encoder. Details about the parameters of this decoder can be found in Table 1. This strategic augmentation aims to enhance the model’s capabilities while maintaining the foundational architecture of BERT encoder, and it also ensures a high degree of comparativeness in our comparative experiments.
@@
Moreover, it is important to note that although the DeBERTa model was trained for the same number of steps as other models, its disentangled attention mechanism introduces a significant overhead comparing with other baseline models and with our BERT+BPDec models. The comparisons can be found in Table 2 This added complexity affects both the pretraining and fine-tuning phases, but ours BERT+BPDec models only have extra ovearhead during pretraining. The implications of this increased overhead are critical, especially when considering the balance between model performance and computational efficiency. Our analysis aims to provide insights into how these different architectures impact training and computation efficiency, while also maintaining or enhancing language processing capabilities.
Models BERT-large BERT-large+BPDec DeBERTa-large
TFLOPs pretraining 1.014 1.182 1.349
Time costs pretraining 100% +17.7% +37.41%
TFLOPs finetuning 1.008 1.176 1.343
Time costs finetuning 100% +0.0% +37.63%
Table 2: Comparison of TFLOPs and Relative Time Costs for Vanilla BERT, BERT+BPDec, and DeBERTa Models. TFLOPs are calculated per data point during pretraining and finetuning, including forward and backward propagation. Time costs are expressed as percentages relative to the Vanilla BERT model (set at 100%).